The First National Bank of Wilson has 640 checking account customers. A recent sample of 60 of these customers showed 23 have a Visa card with the bank.


Construct the 98% confidence interval for the proportion of checking account customers who have a Visa card with the bank. (Use z Distribution Table.) (Round your answers to 3 decimal places.)
Answer #1

OLO E LO Z 24 Joll: 0=60 X=23 , e-g81. for go.. , 2= 2.33 ta Margm of enor = 2 flip) = 0.146 Iener unt=P - E = 0.237 opport c